# io_uring for High-Performance DBMSs: When and How to Use It

### v1

@@ 1. Introduction
Modern PCIe 5.0 hardware, including NVMe SSDs such as the Kioxia CM7-R (2.45M IOPS) and NICs such as the ConnectX-7 (400 Gbit/s), sustains millions of IOPS and hundreds of gigabits per second of throughput, yet conventional I/O interfaces struggle to saturate them (connectx7; DBLP:journals/corr/abs-2506-09226; DBLP:journals/pvldb/HeXLJZYML23; haasWhatModernNVMe2023; DBLP:journals/pvldb/LeisD24). In particular, kernel-based I/O interfaces, as still used by production-grade database systems, incur system-call and context-switch overhead, consuming a significant fraction of CPU cycles without saturating these devices (zhouOLTPLookingGlass2025). These inefficiencies widen the hardware–software gap and make low-overhead I/O mechanisms central to modern database systems.
@@
User-space I/O frameworks such as DPDK, SPDK, and RDMA bypass the kernel and can deliver high performance on dedicated hardware (DBLP:conf/sigmod/LiDSN16; zhouOLTPLookingGlass2025; DBLP:conf/damon/JasnyE0B25). However, operating entirely in user space removes OS abstractions such as file systems and TCP networking, making integration difficult for production databases that rely on them. These stacks also require exclusive control of SSDs or NICs (ButrovichRRLZSP23; FriedCSCGEFB24; KimMBZLPRSSS18), which may conflict with deployments where devices must be shared. Consequently, user-space I/O, despite its advantages, has not seen wide adoption and is used mainly in specialized, tightly controlled environments rather than general-purpose systems (ButrovichRRLZSP23; haasExploitingDirectlyattachedNvme2020).
@@ 3. Efficient Storage I/O with io_uring
Modern NVMe devices can sustain millions of IOPS, yet conventional I/O stacks rarely reach this potential (DBLP:conf/damon/JasnyE0B25; haasWhatModernNVMe2023). To explore if and how io_uring can close this gap, we discuss our three research questions (when to use, how to integrate, and how to tune io_uring) in the context of a buffer-managed storage engine.
@@ 6. Related Work
Although io_uring is a relatively recent addition to the Linux kernel, it has already been studied in various contexts, primarily focusing on storage I/O and its impact on data-intensive systems.
@@
Initial systematic comparisons of io_uring with established interfaces such as libaio and SPDK were conducted using the fio benchmark to evaluate typical storage workloads (didonaUnderstandingModernStorage2022). Ren and Trivedi extended this analysis for Intel Optane SSDs, characterizing microarchitectural behavior, instruction-level overheads, and Linux block I/O scheduler performance (renPerformanceCharacterizationModern2023). In the context of database systems, Haas et al. explored io_uring and other asynchronous I/O mechanisms for NVMe SSD arrays. They identified improvements in throughput and latency associated with various io_uring features (haasExploitingDirectlyattachedNvme2020; haasWhatModernNVMe2023; haasManagingVeryLarge2025).
@@
Several works integrated io_uring into production systems to evaluate application-specific benefits. Chen et al. applied io_uring to Redis, reporting significant reductions in overhead for medium and large payloads (DBLP:conf/cloud2/ChenLLS24). Durner et al. used io_uring to accelerate cloud object storage access, achieving lower latency and higher throughput (durnerExploitingCloudObject2023). However, systematic analyses of network-specific aspects and end-to-end effects on distributed systems remain absent.
@@
From a security perspective, He et al. proposed RingGuard, an eBPF-based framework that monitors and restricts io_uring operations to prevent kernel-level vulnerabilities. It extends eBPF with new io_uring-specific hooks and verifier logic, enforcing safety policies at runtime while maintaining low overhead (heRingGuardGuardIo_uring2023). The most comprehensive exploration of io_uring features to date is the work by Ingimarsson, who integrated basic functionality into RocksDB (ingimarssonExploringPerformanceIo_uring2024). However, detailed evaluations of advanced io_uring features, such as registered buffers for zero-copy operations and linked requests to minimize system calls, are still lacking.

### v2

@@ 1. Introduction
Modern PCIe 5.0 hardware, including NVMe SSDs such as the Kioxia CM7-R (2.45M IOPS) and NICs such as the ConnectX-7 (400 Gbit/s), sustains millions of IOPS and hundreds of gigabits per second of throughput, yet conventional I/O interfaces struggle to saturate them (connectx7; DBLP:journals/corr/abs-2506-09226; DBLP:journals/pvldb/HeXLJZYML23; DBLP:journals/pvldb/HaasL23; DBLP:journals/pvldb/LeisD24). In particular, kernel-based I/O interfaces, as still used by production-grade database systems, incur system-call and context-switch overhead, consuming a significant fraction of CPU cycles without saturating these devices (zhouOLTPLookingGlass2025). These inefficiencies widen the hardware–software gap and make low-overhead I/O mechanisms central to modern database systems.
@@
User-space I/O frameworks such as DPDK, SPDK, and RDMA bypass the kernel and can deliver high performance on dedicated hardware (DBLP:conf/sigmod/LiDSN16; zhouOLTPLookingGlass2025; DBLP:conf/damon/JasnyE0B25). However, operating entirely in user space removes OS abstractions such as file systems and TCP networking, making integration difficult for production databases that rely on them. These stacks also require exclusive control of SSDs or NICs (DBLP:journals/pvldb/ButrovichRRLZSP23; DBLP:conf/nsdi/FriedCSCGEFB24; DBLP:conf/sigcomm/KimMBZLPRSSS18), which may conflict with deployments where devices must be shared. Consequently, user-space I/O, despite its advantages, has not seen wide adoption and is used mainly in specialized, tightly controlled environments rather than general-purpose systems (DBLP:journals/pvldb/ButrovichRRLZSP23; DBLP:conf/cidr/HaasHL20).
@@ 3. Efficient Storage I/O with io_uring
Modern NVMe devices can sustain millions of IOPS, yet conventional I/O stacks rarely reach this potential (DBLP:conf/damon/JasnyE0B25; DBLP:journals/pvldb/HaasL23). To explore if and how io_uring can close this gap, we discuss our three research questions (when to use, how to integrate, and how to tune io_uring) in the context of a buffer-managed storage engine.
@@ 6. Related Work
Although io_uring is a relatively recent addition to the Linux kernel, it has already been studied in various contexts, primarily focusing on storage I/O and its impact on data-intensive systems (DBLP:conf/btw/PestkaP25; DBLP:journals/corr/abs-2411-16254).
@@
Initial systematic comparisons of io_uring with established interfaces such as libaio and SPDK were conducted using the fio benchmark to evaluate typical storage workloads (DBLP:conf/systor/DidonaPIMT22). Ren and Trivedi extended this analysis for Intel Optane SSDs, characterizing microarchitectural behavior, instruction-level overheads, and Linux block I/O scheduler performance (DBLP:conf/eurosys/RenT23). In the context of database systems, Haas et al. explored io_uring and other asynchronous I/O mechanisms for NVMe SSD arrays. They identified improvements in throughput and latency associated with various io_uring features (DBLP:conf/cidr/HaasHL20; DBLP:journals/pvldb/HaasL23; haasManagingVeryLarge2025).
@@
Several works integrated io_uring into production systems to evaluate application-specific benefits. Chen et al. applied io_uring to Redis, reporting significant reductions in overhead for medium and large payloads (DBLP:conf/cloud2/ChenLLS24). Durner et al. used io_uring to accelerate cloud object storage access, achieving lower latency and higher throughput (DBLP:journals/pvldb/DurnerL023). However, systematic analyses of network-specific aspects and end-to-end effects on distributed systems remain absent.
@@
From a security perspective, He et al. proposed RingGuard, an eBPF-based framework that monitors and restricts io_uring operations to prevent kernel-level vulnerabilities. It extends eBPF with new io_uring-specific hooks and verifier logic, enforcing safety policies at runtime while maintaining low overhead (DBLP:conf/sigcomm/HeLZW23). The most comprehensive exploration of io_uring features to date is the work by Ingimarsson, who integrated basic functionality into RocksDB (ingimarssonExploringPerformanceIo_uring2024). However, detailed evaluations of advanced io_uring features, such as registered buffers for zero-copy operations and linked requests to minimize system calls, are still lacking.